Property worth

$249,000.00

can be purchased for 15% down and

monthly

mortgage payments of

$ 1325.00

for

25

years. What effective annual rate of interest is charged?
Answer #1

Using financial calculator
PV=-249000*(1-15%)
PMT=1325
N=12*25
FV=0
CPT I/Y=0.475%

Hence, effective annual rate=(1+0.475%)^12-1=5.851%